OpenStudy (anonymous):

The perimeter of a rectangle is 30 ft. The length is 1 ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting systems. Let x be the length and y be the width. Write the first equation. 2x+2y=______ write the second equation, x=y+______ What is the length____ft. What is the width____ft.

OpenStudy (anonymous):

A/Q, 2x+2y=30...(given) x=y+1 on simplification ,we get, x+y=15 x-y=1 thus solving the above two equations 2x=16 or x=8 y=7 thus, length=8ft and width=7ft :)